OpenStudy (anonymous):

One angle of a quadrilateral has measure a degrees. Find the average of the measures of the other three angles. ( Hint: The sum of the measures of the angles of a quadrilateral is 360 degrees)

OpenStudy (anonymous):

@Dobby1

OpenStudy (dobby1):

hm

OpenStudy (dobby1):

is there answer choices to look at

OpenStudy (anonymous):

no

OpenStudy (dobby1):

hm not sure this is a tricky one

OpenStudy (dobby1):

you did not give the number of degrees for the first angle in the first sentence

OpenStudy (anonymous):

its the variable a and its ( a degrees)

OpenStudy (dobby1):

i am sorry i am pretty darn useles right not i dont want to get you a bad grade so i am just going to find someone who can help

OpenStudy (anonymous):

@mathstudent55

OpenStudy (dobby1):

@amistre64

OpenStudy (mathstudent55):

How do you find the average of 3 numbers?

OpenStudy (anonymous):

hmm i forgot...

OpenStudy (mathstudent55):

Add the three numbers and divide by 3.

OpenStudy (mathstudent55):

This means we need to find the sum of the other three angles. Then we divide that by 3 to find the average.

OpenStudy (mathstudent55):

Notice the hint at the end of the problem.

OpenStudy (mathstudent55):

The sum of the measures of the angles of a quadrilateral is 360 degrees.

OpenStudy (mathstudent55):

You know one angle measures a degrees. All 4 angles together measure 360 degrees. What do the other 3 angles measure together?

OpenStudy (anonymous):

360 degrees/ 3 angles= 120 degrees

OpenStudy (mathstudent55):

No. 1 angle measures a All 4 angles measure 360. The other 3 angles measure 360 - a

OpenStudy (anonymous):

oh so the sum of the other 3 angles is 360-a

OpenStudy (mathstudent55):

 a + m1 + m2 + m3 = 360 m1 + m2 + m3 = 360 - a